# Bilateral Distribution Compression: Reducing Both Data Size and Dimensionality

### v4

@@ Abstract
Existing distribution compression methods reduce dataset size by minimising the Maximum Mean Discrepancy (MMD) between original and compressed sets, but modern datasets are often large in both sample size and dimensionality. We propose Bilateral Distribution Compression (BDC), a two-stage framework that compresses along both axes while preserving the underlying distribution, with overall linear time and memory complexity in dataset size and dimension. Central to BDC is the Decoded MMD (DMMD), which quantifies the discrepancy between the original data and a compressed set decoded from a low-dimensional latent space. BDC proceeds by (i) learning a low-dimensional projection using the Reconstruction MMD (RMMD), and (ii) optimising a latent compressed set with the Encoded MMD (EMMD). We show that this procedure minimises the DMMD, guaranteeing that the compressed set faithfully represents the original distribution. Experiments show that across a variety of scenarios BDC can achieve comparable or superior performance to ambient-space compression at substantially lower cost.
@@ 1 Introduction
Existing distribution compression methods include Kernel Herding (Chen et al., 2010; Bach et al., 2012), Kernel Thinning (Dwivedi and Mackey, 2021b; Shetty et al., 2022; Dwivedi and Mackey, 2021a; Gong et al., 2025), and Gradient Flow approaches (Arbel et al., 2019; Chen et al., 2025). These methods minimise the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) between the empirical distributions of the compressed set $\hat{\mathbb{P}}_{Z}$, and the target dataset $\hat{\mathbb{P}}_{X}$. However, these approaches focus exclusively on reducing the number of observations ($m\ll n$), overlooking the fact that modern datasets are often large both in sample size and in dimensionality. Dimensionality reduction techniques typically focus on preserving pairwise similarities or distances, popular examples include PCA (Maćkiewicz1993PCA; Schölkopf1998KernelPCA), t-SNE (van der Maaten and Hinton, 2008), UMAP (McInnes et al., 2020), and autoencoders (Goodfellow et al., 2016). These methods map data into lower-dimensional spaces ($p\ll d)$ but do not reduce the number of observations.
To address this gap, we propose BDC, a two-stage method that minimises the Decoded Maximum Mean Discrepancy (DMMD), which measures how well a compressed set, once decoded, preserves the original distribution in ambient space. Directly optimising the DMMD is challenging, as it requires simultaneous optimisation of both the decoder and the compressed set. Moreover, a highly flexible decoder risks overfitting by simply memorising the target distribution rather than learning a meaningful representation. To avoid this, we adopt a two-stage procedure:
@@
    In Section 3, we formulate the problem of compressing datasets in both observations ($m\ll n$) and dimensionality ($p\ll d$), and introduce the Decoded MMD (DMMD) as a natural objective for measuring distributional similarity.
@@
    In Section 4, we present empirical results across a range of tasks, datasets, and metrics, showing that BDC is able to match or surpasses the performance of ambient-space distribution compression methods while incurring lower cost.
@@ 2 Preliminaries and Related Work
where $X,X^{\prime}\sim\mathbb{P}_{X}$ and $Y,Y^{\prime}\sim\mathbb{P}_{Y}$. For a broad class of characteristic kernels—including standard choices such as Gaussian, Laplacian, and Matérn kernels (Sriperumbudur et al., 2011)—we have $\text{MMD}(\mathbb{P}_{X},\mathbb{P}_{Y})=0$ if and only if $\mathbb{P}_{X}=\mathbb{P}_{Y}$ (Gretton et al., 2012).
Related Work. Dataset condensation methods typically construct small synthetic datasets optimised for a specific downstream task, most commonly image classification (Sachdeva and McAuley, 2023). The closest method to ours is M3D (Zhang et al., 2024), which performs distribution matching by optimising samples in ambient space using MMD computed in the latent space of randomly initialised encoders. While effective for classification, M3D provides no guarantees of distributional preservation and does not perform dimensionality reduction. Recent neural-based condensation methods (Cazenavette et al., 2023; Moser et al., 2024), including adversarial and diffusion-based approaches, are likewise tailored to image classification and rely on pre-trained generative models to synthesise images from latent codes. These methods do not yield reusable low-dimensional representations and are not applicable beyond domains where suitable generators exist. In contrast, BDC compresses both sample size and data dimension, explicitly targets distributional preservation with theoretical guarantees, and is task- and domain-agnostic. We additionally compare against the state of the art for standard (non-bilateral) distribution compression (Arbel et al., 2019; Chen et al., 2025), and consistently observe improved performance.
@@ 3.1.1 Latent Space Optimisation Objective
To optimise an informative latent space, we first require a suitable loss function. For this purpose, we define the Reconstruction Maximum Mean Discrepancy (RMMD) as
@@
$-\frac{2}{nm}\sum_{i,j=1}^{n}k(\bm{x}_{i},\tilde{\bm{x}}_{j})+\frac{1}{m^{2}}\sum_{i,j=1}^{n}k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j}),$
@@ 3.1.2 Motivating the RMMD
Minimising RMMD with a quadratic kernel identifies the same subspace as PCA. This is intuitive: PCA assumes zero-mean data and identifies directions of maximal variance, while the MMD between two zero-mean distributions under the polynomial kernel is sensitive only to differences in second-order moments (Sriperumbudur et al., 2010). Hence, choosing a characteristic kernel makes RMMD sensitive to all distributional moments, yielding an autoencoder that preserves more distributional information than PCA’s second-order moment matching.
@@ 3.1.3 Latent Space Compression Objective
Given an encoder $\psi$ mapping into the latent space, we require an objective function to optimise the compressed set. For this purpose, we define the Encoded Maximum Mean Discrepancy (EMMD) as
$\text{EMMD}(\hat{\mathbb{P}}_{\psi(X)},\hat{\mathbb{P}}_{Z}):=\frac{1}{n^{2}}\sum_{i,j=1}^{n}h(\psi(\bm{x}_{i}),\psi(\bm{x}_{j}))$ (4)
@@ 2. Step 2:
The overall time and memory cost of the two-stage procedure scales as $\mathcal{O}(nd)$, where $n$ denotes the dataset size and $d$ its dimensionality (see Section C.5). The complexity is therefore linear in both parameters, allowing the algorithm to scale to very large datasets in practice.
As our goal is to minimise the DMMD, the following provides a sufficient condition under which this is achieved by the above method:
@@
Theorem 3.3 provides an ideal-case guarantee: if both RMMD and EMMD are reduced to zero, then DMMD is also exactly zero. Theorem 3.5 instead bounds DMMD by the sum of RMMD and EMMD. In practice, Theorem 3.3 is sufficient to justify the two-stage procedure, while Theorem 3.5 offers an explicit bound on DMMD if required, at the cost of prescribing the latent-space kernel.
@@ 3.2.2 Nonlinear Autoencoders
Neural autoencoders were originally introduced for nonlinear dimensionality reduction (Hinton and Salakhutdinov, 2006). Compared to linear methods, they often achieve superior performance on downstream tasks by exploiting greater flexibility to learn complex nonlinear patterns in data (Fournier and Aloise, 2019). However, excessive flexibility can produce poor latent representations. In the extreme, the network may approximate the identity map, yielding latents that are uninformative for downstream use. Regularisation is therefore essential. Common strategies include enforcing a bottleneck by restricting the size of hidden layers and the final latent dimension, tying encoder and decoder weights, or limiting decoder capacity, each of which reduces the risk of memorisation and encourages the latent space to capture meaningful structure rather than simply copying the input (Goodfellow et al., 2016).
@@
$\qquad+\frac{1}{nd}\sum_{i=1}^{n}\|\bm{x}_{i}-\phi(\psi(\bm{x}_{i}))\|_{2}^{2}.$
See Section A.2.1 for further discussion.
@@ 3.3 Extension to Labelled Data
So far we have assumed that the target dataset is unlabelled. In many applications, however, responses $y\in\mathcal{Y}$ are also available, so that $\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{n}\subset\mathbb{R}^{d}\times\mathcal{Y}$. BDC extends naturally to this supervised setting. Let $l:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ be a kernel on the response space and let $W:\Omega\to\mathcal{Y}$ denote the random variable corresponding to the compressed responses. It is then straightforward to define supervised analogues of our MMD objectives,
$\mathrm{DMMD}\left(\hat{\mathbb{P}}_{X,Y},\,\hat{\mathbb{P}}_{\phi(Z),W}\right),$
$\mathrm{RMMD}\left(\hat{\mathbb{P}}_{X,Y},\,\hat{\mathbb{P}}_{\phi(\psi(X)),Y}\right),$
$\mathrm{EMMD}\left(\hat{\mathbb{P}}_{\psi(X),Y},\,\hat{\mathbb{P}}_{Z,W}\right),$
which employ tensor product kernels (see Section C.3),
@@ 4 Experiments
We evaluate BDC on real and synthetic datasets across both labelled and unlabelled tasks. As ambient-space baselines, we consider: (i) uniform random subsampling (URS); (ii) ambient-space distribution compression via gradient descent (ADC) (Arbel et al., 2019; Broadbent et al., 2025); and (iii) the full dataset (FULL), where computationally feasible. Within BDC, we compare two variants: BDC-L, which employs a linear autoencoder trained on the Stiefel manifold, and BDC-NL, which employs a nonlinear neural autoencoder. For image classification, we additionally compare against M3D, as discussed in the related work.
Across all experiments, the number of observations considered matches or exceeds those typically studied in the distribution compression literature, while the ambient dimensionalities extend substantially beyond standard experimental setups, allowing us to evaluate performance in regimes where both sample size and dimension are large (Chen et al., 2010; Arbel et al., 2019; Dwivedi and Mackey, 2021b). Unless otherwise stated, we use Gaussian kernels defined via the median heuristic (Garreau et al., 2018), and define kernels directly on the learned latent spaces, without employing the pull-back kernel. For further details and additional experiments, including ablation studies on latent dimensionality and kernel choice, see Section D.
@@
Table 1: Overall time comparison (construction of the compressed set plus downstream task fitting) in seconds, across datasets, reported as mean and standard deviation. NA indicates the method cannot be applied to the tesk, OOM indicates computational capacity is exceeded.
@@ 4.1.1 Regression
Swiss-Roll: We first consider a synthetic dataset lying on the Swiss-roll manifold embedded in $\mathbb{R}^{200}$ (Yang and Dunson, 2014). We generate n=20,000n=20{,}000 samples $\bm{x}=(t_{1},t_{2},t_{3})^{\top}$ by
@@
Gaussian noise $\mathcal{N}(\bm{0},I_{3})$ is then added to $\bm{x}$. data are nonlinearly projected to ambient space via a randomly initialised 2-layer feed-forward neural net.
@@
We compress to $m=200$ labelled pairs using the extension from Section 3.3, with latent dimension $p=3$ ($99.99\%$ compression). Figures 3 and 4, and Table 1 show that BDC-NL consistently outperforms ADC across in terms of accuracy and uncertainty on $n_{t}=1{,}000$ test pairs, while being 20× faster. As expected, BDC-L performs poorly due to the misspecification of a linear autoencoder in this nonlinear setting.
@@
Figure 6: Left: test negative log-likelihood on CT-Slice. Right: test continuous ranked probability score on MNIST. BDC-L (red), BDC-NL (green), and ADC (blue) each reported over $10$ runs, URS (grey) over $200$ runs.
@@ 4.1.2 Classification
Figure 7: Test mean squared error on MNIST. BDC-L (red), BDC-NL (green), and ADC (blue) each reported over $10$ runs, URS (grey) over $200$ runs.
Figure 8: Left: test negative log-likelihood on MNIST. Right: test continuous ranked probability score on MNIST. BDC-L (red), BDC-NL (green), and ADC (blue) each reported over $10$ runs, URS (grey) over $200$ runs.
@@ 4.2 Clustering
We construct a synthetic dataset of n=100,000n=100{,}000 two-dimensional points arranged in $10$ clusters of varying shapes, with an additional 2,0002{,}000 points of uniformly distributed noise (see Figure 10). The data are then projected nonlinearly to $d=500$ dimensions using a four-layer decoder with Tanh activations. To recover the clusters, we apply HDBSCAN (Campello et al., 2013), which can detect non-elliptical structures while leaving uncertain points, such as noise, unclustered. We compress to $m=300$ points with latent dimension of $p=2$ (BDC-NL, $99.999\%$ compression) and $p=5$ (BDC-L, $99.997\%$). Ground-truth labels are used only for evaluation where we use an additional $n_{t}=20{,}400$ held-out points. Both BDC-L and BDC-NL are misspecified, with BDC-NL restricted to a one-hidden-layer ReLU autoencoder. Figure 9 and Table 1 show that BDC achieves cluster recovery performance comparable to ADC, while requiring substantially less runtime. Figure 10 also illustrates the original intrinsic data alongside the compressed set constructed by BDC-NL.

### v5

@@ Abstract
Existing distribution compression methods reduce the number of observations in a dataset by minimising the Maximum Mean Discrepancy (MMD) between original and compressed sets, but modern datasets are often large in both sample size and dimensionality. We propose Bilateral Distribution Compression (BDC), a two-stage framework that compresses along both axes while preserving the underlying distribution, with overall linear time and memory complexity in dataset size and dimension. Central to BDC is the Decoded MMD (DMMD), which we introduce to quantify the discrepancy between the original data and a compressed set decoded from a low-dimensional latent space. BDC proceeds by (i) learning a low-dimensional projection using the Reconstruction MMD (RMMD), and (ii) optimising a latent compressed set with the Encoded MMD (EMMD). We show that this procedure minimises the DMMD, guaranteeing that the compressed set faithfully represents the original distribution. Experiments show that BDC can achieve comparable or superior downstream task performance to ambient-space compression at substantially lower cost and with significantly higher rates of compression.
@@ 1 Introduction
Existing distribution compression methods include Kernel Herding (Chen et al., 2010; Bach et al., 2012), Kernel Thinning (Dwivedi and Mackey, 2021; Shetty et al., 2022; Dwivedi and Mackey, 2022; Gong et al., 2025), and Gradient Flow approaches (Arbel et al., 2019; Chen et al., 2025). These methods minimise the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) between the empirical distributions of the compressed set $\hat{\mathbb{P}}_{Z}$, and the target dataset $\hat{\mathbb{P}}_{X}$. However, these approaches focus exclusively on reducing the number of observations ($m\ll n$), overlooking the fact that modern datasets are often large both in sample size and in dimensionality. Dimensionality reduction techniques typically focus on preserving pairwise similarities or distances, popular examples include PCA (Maćkiewicz1993PCA; Schölkopf1998KernelPCA), t-SNE (van der Maaten and Hinton, 2008), UMAP (McInnes et al., 2020), and autoencoders (Goodfellow et al., 2016). These methods map data into lower-dimensional spaces ($p\ll d)$ but do not reduce the number of observations.
To address this gap, we propose BDC, a two-stage method that reduces both data size and dimensionality, by targeting the Decoded Maximum Mean Discrepancy (DMMD), which measures how well a bilaterally compressed set, once decoded, preserves the original distribution in ambient space. Directly optimising the DMMD is challenging, as it requires simultaneous optimisation of both the decoder and the compressed set. Moreover, a highly flexible decoder risks overfitting by simply memorising the target distribution rather than learning a meaningful latent representation. To avoid this, we adopt a two-stage procedure:
@@
    In Section 3, we formulate for the first time the problem of compressing datasets in both observations ($m\ll n$) and dimensionality ($p\ll d$), and introduce the DMMD as a natural objective for measuring distributional similarity.
@@
    In Section 4, we present empirical results across a range of tasks, datasets, and metrics, showing that when manifold structure is present, BDC is able to match or surpasses the performance of ambient-space distribution compression methods while incurring substantially lower cost.
@@ 2 Preliminaries and Related Work
where $X,X^{\prime}\sim\mathbb{P}_{X}$ and $Y,Y^{\prime}\sim\mathbb{P}_{Y}$. For a broad class of characteristic kernels, including standard choices such as Gaussian, Laplacian, and Matérn kernels (Sriperumbudur et al., 2011), we have $\text{MMD}(\mathbb{P}_{X},\mathbb{P}_{Y})=0$ if and only if $\mathbb{P}_{X}=\mathbb{P}_{Y}$ (Gretton et al., 2012).
Related Work. Dataset condensation methods typically construct small synthetic datasets optimised for a specific downstream task, most commonly image classification (Sachdeva and McAuley, 2023). The closest method to ours is M3D (Zhang et al., 2024), which performs distribution matching by optimising samples in ambient space using MMD computed in the latent space of randomly initialised encoders. M3D provides no guarantees of distributional preservation, does not perform dimensionality reduction and is designed only for classification tasks. Recent neural-based condensation methods (Cazenavette et al., 2023; Moser et al., 2024), including adversarial and diffusion-based approaches, are likewise tailored to image classification and rely on pre-trained generative models to synthesise images from latent codes. These methods do not yield reusable low-dimensional representations and are not applicable beyond domains where suitable generators exist. In contrast, BDC compresses both sample size and data dimension, explicitly targets distributional preservation with theoretical guarantees, and is task- and domain-agnostic.
@@ 3.1.1 Latent Space Optimisation Objective
To achieve distribution-preserving dimensionality reduction ($p\ll d$), we must optimise an informative latent space using a suitable loss function. For this purpose, we introduce the Reconstruction Maximum Mean Discrepancy (RMMD) as
@@
$-\frac{2}{nm}\sum_{i,j=1}^{n,m}k(\bm{x}_{i},\tilde{\bm{x}}_{j})+\frac{1}{m^{2}}\sum_{i,j=1}^{m}k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j}),$
@@ 3.1.2 Motivating the RMMD
Minimising the RMMD with a quadratic kernel identifies the same subspace as PCA. This is intuitive: PCA assumes zero-mean data and identifies directions of maximal variance, while the MMD between two zero-mean distributions under the polynomial kernel is sensitive only to differences in second-order moments (Sriperumbudur et al., 2010). Hence, choosing a characteristic kernel makes RMMD sensitive to all distributional moments, yielding an autoencoder that preserves more distributional information than PCA’s second-order moment matching.
@@ 3.1.3 Latent Space Compression Objective
Given an encoder $\psi$ mapping into a low-dimensional latent space ($p\ll d$), we require an objective function to optimise a compressed set ($m\ll n$) that preserves the encoded data distribution. For this purpose, we define the Encoded Maximum Mean Discrepancy (EMMD) as
$\text{EMMD}^{2}(\hat{\mathbb{P}}_{\psi(X)},\hat{\mathbb{P}}_{Z}):=\frac{1}{n^{2}}\sum_{i,j=1}^{n}h(\psi(\bm{x}_{i}),\psi(\bm{x}_{j}))$ (4)
@@ 2. Step 2:
The overall time and memory cost of the two-stage procedure scales as $\mathcal{O}(nd)$, where $n$ denotes the dataset size and $d$ its dimensionality (see Section C.5). The complexity is therefore linear in both parameters, allowing the algorithm to scale to large datasets.
As our goal is to minimise the DMMD, the following theorem provides a sufficient condition under which this is achieved by the above method:
@@
Theorem 3.3 justifies the two-stage procedure, if both RMMD and EMMD are optimised to zero, then DMMD is also zero. Theorem 3.5 offers an explicit bound on DMMD if required, at the cost of prescribing the latent-space kernel.
@@ 3.2.2 Nonlinear Autoencoders
Neural autoencoders were originally introduced for nonlinear dimensionality reduction (Hinton and Salakhutdinov, 2006). Compared to linear methods, they may achieve superior performance on downstream tasks by exploiting greater flexibility to learn complex nonlinear patterns in data (Fournier and Aloise, 2019). However, excessive flexibility can produce poor latent representations. In the extreme, the network may approximate the identity map, yielding latents that are uninformative for downstream use. Regularisation is therefore essential. Common strategies include enforcing a bottleneck by restricting the size of hidden layers and the final latent dimension, tying encoder and decoder weights, or limiting decoder capacity, each of which reduces the risk of memorisation and encourages the latent space to capture meaningful structure rather than simply copying the input (Goodfellow et al., 2016).
@@
$\qquad+\frac{1}{nd}\sum_{i=1}^{n}\|\bm{x}_{i}-\phi(\psi(\bm{x}_{i}))\|_{2}^{2}$
which combines their strengths. See Section A.2.1 for further discussion, including a toy example illustrating why RMMD alone is insufficient when using highly flexible autoencoders. [^2] In Section D.1.9, we introduce a weighting parameter into Equation 5, and show that performance is robust to it provided the MSRE term is not weighted arbitrarily small.
@@ 3.3 Extension to Labelled Data
So far we have assumed that the target dataset is unlabelled. In many applications, however, responses $y\in\mathcal{Y}$ are also available, so that $\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{n}\subset\mathbb{R}^{d}\times\mathcal{Y}$. BDC extends naturally to this setting. Let $l:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ be a kernel on the response space and let $W:\Omega\to\mathcal{Y}$ denote the r.v. of the compressed responses. It is then straightforward to define supervised analogues of our MMD objectives,
$\mathrm{DMMD}\left(\hat{\mathbb{P}}_{X,Y},\,\hat{\mathbb{P}}_{\phi(Z),W}\right),\;\mathrm{EMMD}\left(\hat{\mathbb{P}}_{\psi(X),Y},\,\hat{\mathbb{P}}_{Z,W}\right),$
$\qquad\qquad\qquad\mathrm{RMMD}\left(\hat{\mathbb{P}}_{X,Y},\,\hat{\mathbb{P}}_{\phi(\psi(X)),Y}\right),$
which employ tensor product kernels to target the joint distribution (see Section C.3 for details),
@@ 4 Experiments
We evaluate BDC on real and synthetic datasets across both labelled and unlabelled tasks. As ambient-space baselines, we consider: (i) uniform random subsampling (URS); (ii) ambient-space distribution compression via gradient descent (ADC) (Arbel et al., 2019); and (iii) the full dataset (FULL), where computationally feasible. Within BDC, we compare two variants: BDC-L, which employs a linear autoencoder trained on the Stiefel manifold, and BDC-NL, which employs a nonlinear neural autoencoder. For image classification, we additionally compare against M3D, as discussed in the related work.
Across all experiments, the number of observations considered matches or exceeds those typically studied in the distribution compression literature (Chen et al., 2010; Arbel et al., 2019; Dwivedi and Mackey, 2021), moreover, the ambient dimensionalities extend substantially beyond standard experimental setups, allowing us to evaluate performance in regimes where both sample size and dimension are large. Unless otherwise stated, we use Gaussian kernels defined via the median heuristic (Garreau et al., 2018), and define kernels directly on the learned latent spaces, without employing the pull-back kernel. For further details and additional experiments, including ablation studies on the latent dimensionality $p$, the number of compressed set points $m$, and the choice of kernels, see Section D.
@@
Table 1: Overall time comparison (construction of the compressed set plus downstream task fitting) in seconds, across datasets, reported as mean and standard deviation. NA indicates the method cannot be applied to the task, OOM indicates computational capacity is exceeded.
@@ 4.1.1 Regression
Swiss-Roll: We first consider a synthetic dataset lying on the Swiss-roll manifold embedded in $\mathbb{R}^{200}$ (Yang and Dunson, 2016). We generate n=20,000n=20{,}000 samples $\bm{x}=(t_{1},t_{2},t_{3})^{\top}$ by
@@
Gaussian noise $\mathcal{N}(\bm{0},I_{3})$ is then added to $\bm{x}$. Data are nonlinearly projected to 200 dimensions via a randomly initialised 2-layer feed-forward neural net. Responses are defined as
@@
We compress to $m=200$ labelled pairs using the extension from Section 3.3, with latent dimension $p=3$ ($99.99\%$ compression). Figures 3 and 4, and Table 1 show that BDC-NL consistently outperforms ADC across in terms of accuracy and uncertainty on $n_{t}=1{,}000$ test pairs, while being $20\times$ faster and achieving almost $70\times$ more compression. As expected, BDC-L performs poorly due to the misspecification of a linear autoencoder in this nonlinear setting.
@@
Figure 6: Left: test negative log-likelihood on CT-Slice. Right: test continuous ranked probability score on CT-Slice. BDC-L (red), BDC-NL (green), and ADC (blue) each reported over $10$ runs, URS (grey) over $200$ runs.
@@ 4.1.2 Classification
Figure 7: Test classification accuracy on MNIST. BDC-L (red), BDC-NL (green), and ADC (blue) each reported over $10$ runs, URS (grey) over $200$ runs.
Figure 8: Left: test negative log-likelihood on MNIST. Right: test multi-class Brier score on MNIST. BDC-L (red), BDC-NL (green), and ADC (blue) each reported over $10$ runs, URS (grey) over $200$ runs.
@@ 4.2 Clustering
We construct a synthetic dataset of n=100,000n=100{,}000 two-dimensional points arranged in $10$ clusters of varying shapes, with an additional 2,0002{,}000 points of uniformly distributed noise (see Figure 10). The data are then projected nonlinearly to $d=500$ dimensions using a four-layer decoder with Tanh activations. To recover the clusters, we apply HDBSCAN (Campello et al., 2013), which can detect non-elliptical structures while leaving uncertain points, such as noise, unclustered. We compress to $m=300$ points with latent dimension of $p=2$ (BDC-NL, $99.999\%$ compression) and $p=5$ (BDC-L, $99.997\%$ compression). Ground-truth labels are used only for evaluation where we use an additional $n_{t}=20{,}400$ held-out points. Both BDC-L and BDC-NL are misspecified, with BDC-NL restricted to a one-hidden-layer ReLU autoencoder. Figure 9 and Table 1 show that BDC achieves cluster recovery performance comparable to ADC, while requiring substantially less runtime. Figure 10 also illustrates the original intrinsic data alongside the compressed set constructed by BDC-NL.